# A Coarse-to-Fine Approach to Multi-Modality 3D Occupancy Grounding

### v1

@@ IV-B Talk2Occ Baselines
[…]
@@ IV-C GroundingOcc
[…]
@@ V-A Implementation Details
[…]
For the image backbone, we use ResNet101-DCN [54], providing multi-scale features at resolutions of 1/81/81 / 8, 1/161/161 / 16, 1/321/321 / 32, and 1/641/641 / 64. The text encoder is a pre-trained RoBERTa-base model [55]. Depth maps are estimated from the $P_{3}$ feature map at 1/81/81 / 8 resolution. The voxel encoder consists of 333 layers with 444 sampling points per voxel query. We fuse data from 444 consecutive frames (0.50.50.5s apart) with single-frame LiDAR point clouds to generate fused voxel embeddings. The occupancy head uses two MLP layers with 128128128 hidden units and softplus activation, while the grounding head has 333 layers with grounding queries of dimension 256256256.
@@
We train both the baselines and GroundingOcc model on 4 NVIDIA RTX 4090 GPUs, with a batch size of 1 per GPU. The AdamW optimizer is adopted with an initial learning rate of $2\times 10^{-4}$, cosine annealing schedule, and weight decay of 0.010.010.01. Data augmentation techniques such as image cropping, color distortion, and GridMask [56] are applied. Input images are cropped and resized to 320×800320$\times$ 800320 × 800 for training due to GPU memory constraints.

### v2

@@ IV-B Talk2Occ Baselines
As depicted at the bottom of Fig. 3, we first employ an image encoder to extract multi-view features $F^{mv}$. Inspired by the querying paradigm in recent works [43, 42], we introduce a 3D voxel query mechanism($Q_{vq}$). This mechanism extracts features from the multi-view features $F^{mv}$ and prompt embeddings $F^{\mathcal{T}}\in\mathbb{R}^{L\times d}$. This process outputs camera voxel features $F^{\mathcal{C}}$, maintaining the same volumetric size as $F^{\mathcal{L}}$. Following the LiDAR-based baseline, we apply the voxel decoder, occupancy head, and grounding head to generate the final occupancy grounding results, denoted as $O^{\mathcal{OG-C}}\in\mathbb{R}^{H\times W\times Z}$.
@@ IV-C GroundingOcc
where FFN, MHCA, and MHSA refer to the feed-forward network, multi-head cross-attention, and multi-head self-attention layers, respectively. Following [47], convolutional layers predict the 2D grounding outputs, which include the 2D bounding box coordinates(l,r,t,b)(l,r,t,b), center $(x_{2D},y_{2D})$, and a centerness measure $d$.
@@ V-A Implementation Details
Our implementation is based on the nuScenes dataset [16], like Occ3D-nuScenes [15]. The point cloud range is $\left[-40.0\,\text{m},40.0\,\text{m}\right]$ for the $x$ and $y$ axes, and $\left[-1.0\,\text{m},5.4\,\text{m}\right]$ for the $z$ axis. We use a voxel grid size of $(0.4\,\text{m},0.4\,\text{m},0.4\,\text{m})$ for loss supervision and an initial voxel query resolution of $50\times 50\times 16$, with embedding dimension $D=256$. After upsampling, the resolution becomes $200\times 200\times 32$ with $D^{\prime}=64$.
For the image backbone, we use ResNet101-DCN [54], providing multi-scale features at resolutions of $1/8$, $1/16$, $1/32$, and $1/64$. The text encoder is a pre-trained RoBERTa-base model [55]. Depth maps are estimated from the $P_{3}$ feature map at $1/8$ resolution. The voxel encoder consists of $3$ layers with $4$ sampling points per voxel query. We fuse data from $4$ consecutive frames ($0.5$s apart) with single-frame LiDAR point clouds to generate fused voxel embeddings. The occupancy head uses two MLP layers with $128$ hidden units and softplus activation, while the grounding head has $3$ layers with grounding queries of dimension $256$.
@@
We train both the baselines and GroundingOcc model on 4 NVIDIA RTX 4090 GPUs, with a batch size of 1 per GPU. The AdamW optimizer is adopted with an initial learning rate of $2\times 10^{-4}$, cosine annealing schedule, and weight decay of $0.01$. Data augmentation techniques such as image cropping, color distortion, and GridMask [56] are applied. Input images are cropped and resized to $320\times 800$ for training due to GPU memory constraints.